OpenStudy (anonymous):

Solve the word problem. Paula has 24 coins in her drawer. Seven of the coins are nickels and the rest are pennies. What is the total value of the coins in her drawer? A. 7¢ B. 35¢ C. 52¢ D. 120¢

OpenStudy (anonymous):

do 26 -7 = 19. Now solve 7(5) + 19 = ?

OpenStudy (anonymous):

54

OpenStudy (anonymous):

I'm sorry, 24-7 = 17, so you actually have to do 7(5) + 17